OpenStudy (anonymous):

3/10+ 7/8= Put that in simplest form, would that be 10/18 and divide each by 2 and get 5/9?

OpenStudy (mathstudent55):

You need a common denominator for the two fractions.

OpenStudy (amistre64):

3/10 + 7/8 = n 3/5 + 7/4 = 2n 3 + 7(5)/4 = 2(5)n 3(4) + 7(5) = 2(5)(4)n 12 + 35 = 40n 47 = 40n 47/40 = n

OpenStudy (anonymous):

I don't get it.

OpenStudy (mathstudent55):